# Extreme-mass-ratio inspirals in relativistic accretion discs

### v1

@@ I Introduction
Dense astrophysical environments at the centres of galaxies—such as nuclear clusters [1, 2], the discs of active galactic nuclei (AGN) [3], and dark-matter overdensities [4, 5, 6]—have assumed growing interest in gravitational-wave (GW) astronomy. It is now well established that these structures can significantly influence the formation, evolution, and GW signature of compact binaries across the entire GW spectrum [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 17]. This increased activity is partially driven by the recognition that relativistic effects are crucial to accurately model the interaction between binaries and their astrophysical surroundings [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
@@ II Executive Summary
As we were finalizing this work, we became aware of independent efforts [55] pursuing the same problem through a (semi-)analytical Hamiltonian approach, improving on Hirata’s original formulation. The methodologies differ: our work focuses on numerically evaluating small multipoles and extrapolating to large ones, whereas Ref. [55] derives analytical expressions at large multipoles and extrapolates downward. Our findings are consistent, reinforcing the robustness of these results.
@@ III.1 Location of the resonances
As the secondary moves inward and $r_{s}$ decreases, this pattern reverses: the inner Lindblad resonances move closer to the secondary than the outer ones, and the net disc torque may become outward. If GW emission dominates the evolution, this slows down the inspiral with respect to vacuum. If migration torques compete with or dominate over GW emission, the inspiral can be stalled, forming a migration trap for the EMRI. The critical radius where the inner and outer resonance locations switch proximity is spin-dependent—it occurs at smaller $r_{s}$ for higher $a$. Specifically, for a non-spinning primary, $a=0$, this transition happens at $r_{s}\sim 18M$, while for a rapidly spinning one, $a=0.9M$, it is shifted inward to approximately $r_{s}\sim 10M$. This feature is also present in the analysis of Ref. [55].
@@ III.2.3 Results
Since the larger the $m$, the closer the resonant radius is to the secondary, the amplitude of the torques grows with $m$. In Fig. 4, we present this growth for a representative orbital radius of the companion and primary BH spin. We find that the growth of the inner and outer Lindblad torques with $m$ for fixed secondary radius is quadratic already at small $m$, and the total normalised Lindblad torque (i.e., their sum) is approximately linear (Ref. [55] presents an analytical derivation of this scaling in the large $m$ limit). We verified this happens for any spin and even at larger orbital radii. This quadratic scaling provides a substantial computational advantage in evaluating the total Lindblad torque at a given radius (summed over $m$). Rather than computing high-$m$ modes, which require prohibitively many $\ell$-modes, it suffices to calculate the torques at small $m$, perform a quadratic fit, and extrapolate to larger $m$. Without this simplification, obtaining the total torques would be numerically infeasible, as solving the Teukolsky equation at large $\ell$ is challenging.
@@ IV.1 Pressure cut-off
This suppression produces a cut-off function for the torque that combines polynomial and exponential decay. In particular, we adopt the following mode-by-mode prescription for the pressure cutoff [^3] Alternative, one can introduce a cut-off at $\xi=1$, as done in Ref. [55].
@@ IV.3 Results on total torques
The results are displayed in Fig. 6. A key feature, also highlighted in the recent work [55], is that relativistic torques in relativistic discs can differ by orders of magnitude from their Newtonian counterparts. In our Fig. 6, part of this variation arises from the normalisation of the disc profiles in Fig. 5: while these profiles coincide in the Newtonian limit ($r\to\infty$), they differ substantially in the strong-field regime. This discrepancy is not only due to differences in surface densities but also in scale heights, which determine the pressure cut-off and thus the number of contributing Lindblad resonances. Choosing an alternative normalisation, e.g., matching both surface densities and aspect ratios at a smaller radius (such as the ISCO, $r=10M$), still yields large differences in torque amplitudes at other radii.
@@ V Discussion
As one of the first investigations of relativistic Lindblad torques on EMRIs—together with [55], to which our model should be quantitatively compared in the future—our work has some limitations:
@@ VI Acknowledgements
We thank Abhishek Hegade K. R., Nicolás Yunes, and Conor Dyson for fruitful discussions and for generously sharing their results prior to publication. We also thank Zachary Nasipak for his guidance in the use of the package pybhpt, which was essential for efficiently computing the relativistic Lindblad torques. We thank Diogo Ribeiro for valuable discussions on numerical implementations. We thank Thomas Spieksma and Conor Dyson for carefully reading the manuscript and providing valuable feedback. RV gratefully acknowledges the support of the Dutch Research Council (NWO) through an Open Competition Domain Science-M grant, project number OCENW.M.21.375. AS acknowledges support from the STFC Consolidated Grant no. ST/V005596/1.

### v2

@@ I Introduction
Dense astrophysical environments at the centres of galaxies—such as nuclear clusters [1, 2], the discs of active galactic nuclei (AGN) [3], and dark-matter overdensities [4, 5, 6]—have assumed growing interest in gravitational-wave (GW) astronomy. It is now well established that these structures can significantly influence the formation, evolution, and GW signature of compact binaries across the entire GW spectrum [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 17, 21]. This increased activity is partially driven by the recognition that relativistic effects are crucial to accurately model the interaction between binaries and their astrophysical surroundings [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
@@ II Executive Summary
As we were finalizing this work, we became aware of independent efforts [56, 57] pursuing the same problem through a (semi-)analytical Hamiltonian approach, improving on Hirata’s original formulation. The methodologies differ: our work focuses on numerically evaluating small multipoles and extrapolating to large ones, whereas Refs. [56, 57] derive analytical expressions at large multipoles and extrapolates downward. Our findings are consistent, reinforcing the robustness of these results.
@@ III.1 Location of the resonances
As the secondary moves inward and $r_{s}$ decreases, this pattern reverses: the inner Lindblad resonances move closer to the secondary than the outer ones, and the net disc torque may become outward. If GW emission dominates the evolution, this slows down the inspiral with respect to vacuum. If migration torques compete with or dominate over GW emission, the inspiral can be stalled, forming a migration trap for the EMRI. The critical radius where the inner and outer resonance locations switch proximity is spin-dependent—it occurs at smaller $r_{s}$ for higher $a$. Specifically, for a non-spinning primary, $a=0$, this transition happens at $r_{s}\sim 18M$, while for a rapidly spinning one, $a=0.9M$, it is shifted inward to approximately $r_{s}\sim 10M$. This feature is also present in the analysis of Refs. [56, 57].
@@ III.2.3 Results
Since the larger the $m$, the closer the resonant radius is to the secondary, the amplitude of the torques grows with $m$. In Fig. 4, we present this growth for a representative orbital radius of the companion and primary BH spin. We find that the growth of the inner and outer Lindblad torques with $m$ for fixed secondary radius is quadratic already at small $m$, and the total normalised Lindblad torque (i.e., their sum) is approximately linear (Refs. [56, 57] presents an analytical derivation of this scaling in the large $m$ limit). We verified this happens for any spin and even at larger orbital radii. This quadratic scaling provides a substantial computational advantage in evaluating the total Lindblad torque at a given radius (summed over $m$). Rather than computing high-$m$ modes, which require prohibitively many $\ell$-modes, it suffices to calculate the torques at small $m$, perform a quadratic fit, and extrapolate to larger $m$. Without this simplification, obtaining the total torques would be numerically infeasible, as solving the Teukolsky equation at large $\ell$ is challenging.
@@ IV.1 Pressure cut-off
This suppression produces a cut-off function for the torque that combines polynomial and exponential decay. In particular, we adopt the following mode-by-mode prescription for the pressure cutoff [^3] Alternative, one can introduce a cut-off at $\xi=1$, as done in Refs. [56, 57].
@@ IV.3 Results on total torques
The results are displayed in Fig. 6. A key feature, also highlighted in the recent works [56, 57], is that relativistic torques in relativistic discs can differ by orders of magnitude from their Newtonian counterparts. In our Fig. 6, part of this variation arises from the normalisation of the disc profiles in Fig. 5: while these profiles coincide in the Newtonian limit ($r\to\infty$), they differ substantially in the strong-field regime. This discrepancy is not only due to differences in surface densities but also in scale heights, which determine the pressure cut-off and thus the number of contributing Lindblad resonances. Choosing an alternative normalisation, e.g., matching both surface densities and aspect ratios at a smaller radius (such as the ISCO, $r=10M$), still yields large differences in torque amplitudes at other radii.
@@ V Discussion
As one of the first investigations of relativistic Lindblad torques on EMRIs—together with [56, 57], to which our model should be quantitatively compared in the future—our work has some limitations:
@@ VI Acknowledgements
We thank Abhishek Hegade K. R., Nicolás Yunes, and Conor Dyson for fruitful discussions and for generously sharing their results prior to publication. We also thank Zachary Nasipak for his guidance in the use of the package pybhpt, which was essential for efficiently computing the relativistic Lindblad torques. We thank Diogo Ribeiro for valuable discussions on numerical implementations. We thank Thomas Spieksma and Conor Dyson for carefully reading the manuscript and providing valuable feedback. RV gratefully acknowledges the support of the Dutch Research Council (NWO) through an Open Competition Domain Science-M grant, project number OCENW.M.21.375. AS acknowledges support from the STFC Consolidated Grant no. ST/V005596/1. LS acknowledges support from the UKRI Horizon guarantee funding (project no. EP/Y023706/1). LS is also supported by a University of Nottingham Anne McLaren Fellowship. LS would like to thank the TAPIR group at Caltech for their kind hospitality during the initial stages of this work.